For the FY 2016, Alpha Company's balance sheet included the following current items: cash $40,000, accounts receivable $160,000, inventories $100,000, prepaid expenses $20,000, accounts payable $90,000, and accrued expenses $70,000. Use this information to determine the: (Round & enter your answers to one decimal place for non-dollar ratios and enter the value. For dollar ratio enter as whole dollars only.) 1. Quick Ratio 2. Current Ratio 3. Working Capital
Answer #1
Alpha Company's
Current Ratio
Current Assets/Current Liabilities
Current Assets
Cash $                     40,000.00
Accounts Receivable $                  1,60,000.00
Inventory $                  1,00,000.00
Prepaid Expenses $                     20,000.00
Total Current Assets=(A) $                  3,20,000.00

Current Liabilities

Accounts Payable $                     90,000.00
Accrued Expenses $                     70,000.00
Total Current Liabilities=(B) $                  1,60,000.00
Current Ratio=(A)/(B) (2:1)
Quick Ratio
Quick Assets/Current Liabilities
Quick Assets
Cash $                     40,000.00
Accounts Receivable $                  1,60,000.00
Total Quick Assets=(A) $                  2,00,000.00
Current Liabilities
Accounts Payable $                     90,000.00
Accrued Expenses $                     70,000.00
Total Current Liabilities=(B) $                  1,60,000.00
Quick Ratio=(A)/(B) (1.25:1)
Working Capital=Current Assets-Current Liabilities
Current Assets=(A) $                  3,20,000.00
Current Liabilities=(B) $                  1,60,000.00
Working Capital=(A)-(B) $                  1,60,000.00